# Learning 3D Representations from Procedural 3D Programs

### v1

@@ Abstract
Remarkably, despite lacking semantic content, the 3D representations learned from this synthesized dataset perform on par with state-of-the-art representations learned from semantically recognizable 3D models (e.g., airplanes) across various downstream 3D tasks, including shape classification, part segmentation, and masked point cloud completion. Our analysis further suggests that current self-supervised learning methods primarily capture geometric structures rather than high-level semantics.
@@ 1 Introduction
Self-supervised learning (SSL) aims at learning representations from unlabeled data that can transfer effectively to various downstream tasks. Inspired by the success of SSL in language [8] and image representation learning  [10, 11], SSLs for 3D point cloud understanding have recently gained considerable interest [27, 42, 19]. For instance, Point-MAE [19] introduces the masked autoencoding scheme into point cloud representation learning, showing substantial improvements in various 3D shape understanding tasks (e.g., shape classification and segmentation).
However, unlike language and image data, which are abundantly available, 3D assets are less accessible due to the expertise required for creating 3D shapes using specialized software (e.g., Blender) or professional 3D scanning equipment. This scarcity of 3D shapes limits the scalability of existing representation learning methods. While recent efforts aim to expand 3D object datasets [6, 7], challenges unique to 3D data collection — such as copyright issues and format diversity — remain unresolved.
To address these challenges, we explore learning point cloud representations from solely synthetic data generated via procedural 3D programs [34, 37, 38] — examples are shown in Fig. LABEL:fig:teaserb. Our data generation pipeline begins with sampling shapes from a set of simple 3D primitives (e.g., cubes, cylinders, and spheres). These primitives undergo affine transformations (e.g., scaling, translation, and rotation) and are combined to create diverse geometries. We then augment the composed shapes using predefined operations (e.g., Boolean operations) to further enhance topological diversity, and uniformly sample 3D surface points from these shapes for point cloud representation learning. We generate a total of 150K synthetic 3D point clouds, costing around 600 CPU hours, though this pipeline can theoretically produce an unlimited number of 3D shapes without copyright concerns. This leads us to two key questions:
@@
To answer these questions, we leverage Point-MAE to learn representations from purely synthetic 3D shapes. We call this framework Point-MAE-Zero to emphasize that it does not use any human-made 3D shapes. We evaluate Point-MAE-Zero and conduct extensive comparisons with Point-MAE pretrained on ShapeNet (referred to as Point-MAE-SN) across several downstream tasks, including shape classification, part segmentation, and masked point cloud reconstruction. Our key findings are as follows:
  • Surprisingly, despite the lack of semantics in the synthetic dataset, Point-MAE-Zero performs comparably to Point-MAE-SN across various 3D tasks, as shown in Fig. LABEL:fig:teaserc;
  • Point-MAE-Zero and Point-MAE-SN exhibit similar behavior patterns. Both models can reconstruct masked point clouds regardless of their training domain — Point-MAE-Zero accurately predicts missing points in ShapeNet shapes (e.g., airplanes and chairs) as shown in Fig. LABEL:fig:teaserd, while Point-MAE-SN can reconstructs synthetic shapes, as seen in Fig. 2. The two models also share structural similarities in their latent spaces, illustrated in Fig. 6;
  • Point-MAE-Zero demonstrates improved performance with increased geometric diversity and dataset size in the synthetic dataset, as illustrated in Tab. 3 and Fig. 4.
Our work is inspired by and builds upon a recent study that successfully trained large 3D reconstruction models exclusively on procedurally generated shapes [34]. Our exploration is also closely related to prior efforts that learn image representations from procedural programs [2, 1]. Concurrent to our work, Yu et al. [43] demonstrated that procedurally generated videos can perform as effectively as natural videos for self-supervised video representation learning.
@@ 2 Related Work
Figure 1: Point-MAE-Zero. (a) Our synthetic 3D point clouds are generated by sampling, compositing, and augmenting simple primitives with procedural 3D programs [34]. (b) We use Point-MAE [42] as our pretraining framework to learn 3D representation from synthetic 3D shapes, dubbed Point-MAE-Zero where “Zero” underscores that we do not use any human-made 3D shapes. (c) We evaluate Point-MAE-Zero in various 3D shape understanding tasks.
3D Object Datasets.
@@
Large-scale datasets are essential in advancing 3D deep learning. Unlike 2D images or videos, building and annotating 3D models requires expertise with professional 3D software or scanning equipment, making the process more costly and time-intensive. Despite these challenges, significant efforts have been made to curate extensive 3D shape datasets [3, 32, 22, 23, 5, 9, 18, 29, 6]. For example, ShapeNet provides 3 million CAD models, with 51K of them being clean, high-quality models. More recently, Objaverse-XL [7] expanded the 3D dataset to 10.2 million models, though scaling up has introduced challenges, including increased noise, format diversity, and unresolved copyright and legal issues. In contrast, we explore procedural 3D programs that generate 3D shapes from simple primitives. This synthetic approach can theoretically produce an unlimited number of 3D shapes without licensing concerns.
Supervised Learning for 3D Point Clouds.
@@
Unlike 2D images, 3D point clouds are inherently unordered and have an irregular structure, leading to extensive research on specialized neural network architectures [20, 21, 28, 44, 30, 31, 36, 13]. For instance, PointNet [20] introduces permutation-invariant operators and pooling layers to effectively aggregate features across 3D points. PointNet++ [21] builds on this by incorporating a hierarchical spatial structure to capture more localized geometric information. DGCNN [28] adopts a graph-based approach, constructing a graph from input point clouds and applying a graph CNN to aggregate features from unordered 3D points. More recent models, such as the Point Transformer [44, 30, 31], utilize a modified transformer architecture tailored specifically for 3D point cloud processing. In our work, we employ a standard transformer architecture, aligning with recent trends in self-supervised learning [11, 19, 42].
Self-supervised Learning for Point Clouds.
@@
Self-supervised learning (SSL) for 3D representations aims to learn features that transfer well across diverse point cloud tasks, such as shape classification, object detection, and part segmentation. Recent SSL approaches for point clouds generally fall into two categories: contrastive learning [10, 35, 4] and masked autoencoding [27, 42, 19]. PointContrast [35] and DepthContrast [4] use an instance discrimination task [10] to learn 3D representations. OcCo [27] learns point cloud representations by reconstructing the original point cloud from occluded views, while IAE [39] trains an autoencoder to recover implicit features from point cloud inputs. Inspired by the success of masked autoencoding in vision [11] and language representation learning [8], Point-BERT [42] and Point-MAE [19] learn representations by predicting masked regions, using transformers as the underlying architecture. In this work, we adopt Point-MAE as our self-supervised learning method due to its state-of-the-art performance in various 3D understanding tasks.
Learning from Synthetic Data
@@
Synthetic data has become popular in computer vision, especially in scenarios where ground-truth annotations are difficult to obtain or where privacy and copyright issues arise. State-of-the-art performance in mid-level or 3D vision tasks is often achieved through training on synthetic data, including tasks like optical flow [24], depth estimation [40], dense tracking [12], relighting [37], novel view synthesis [38], and material estimation [14]. Procedurally generated synthetic data has also been explored for self-supervised representation learning in images [2, 1] and videos [43], and more recently for multi-view feed-forward 3D reconstruction [34]. Concurrent with our study, Yu et al. [43] investigates self-supervised video representation learning from procedurally generated images and videos. In this work, we explore self-supervised representation learning for point clouds, using synthetic 3D shapes generated by procedural 3D programs.
@@ 3.1 Procedural 3D programs
There is a line of works that synthesize 3D shapes using procedural 3D programs for vision tasks, such as novel view synthesis [38], relighting [37], and material estimation [14]. We closely follow the recent work [34] that collects procedurally synthesized 3D shapes for training sparse-view 3D reconstruction models, while we tackle a different task in this wrok — self-supervised 3D representation learning from purely synthetic 3D datasets. Fig. 1a illustrates our data collection pipeline:
  • (1). The procedural 3D program randomly samples $\mathbf{K}$ shapes from a set of simple primitive shapes, including cubes, spheres, cylinders, cones, and tori. These primitives are transformed with affine transformations and combined;
  • (2). To enrich geometric diversity, the composed shapes undergo various predefined manipulations, such as boolean difference and wireframe conversion. For more details on shape augmentation, refer to Xie et al. [34];
  • (3). We uniformly sample $\mathbf{N}$ surface points from the synthesized shapes, which serve as input to the model for representation learning, as illustrated in Fig. 1b.
We explore different configurations for shape generation, such as varying the maximum number of primitives sampled in step (1) and including or excluding shape augmentation operations in step (2). By default, each dataset configuration includes 150K generated shapes, with $\mathbf{N}=8192$ surface points per shape. A detailed analysis of the impact of dataset size and shape complexity on 3D representation learning is provided in Sec. 4.
3.2 Point-MAE-Zero
@@
As depicted in Fig. 1b, we use Point-MAE[11], a state-of-the-art method for point cloud representation learning, to train on synthesized 3D shapes. Point-MAE learns 3D representations through a masked autoencoding scheme [8, 11, 42]. Specifically, it segments the input point cloud into irregular point patches and randomly masks a large proportion of them (60% by default). Shape representations are then learned using a transformer-based encoder-decoder by reconstructing the masked point patches.
The reconstruction loss is computed as the $L_{2}$ Chamfer Distance between the predicted point patches $P_{\text{pre}}$ and the ground-truth patches $P_{\text{gt}}$:
$L=\frac{1}{|P_{\text{pre}}|}\sum_{a\in P_{\text{pre}}}\min_{b\in P_{\text{gt}}}\|a-b\|_{2}^{2}+\frac{1}{|P_{\text{gt}}|}\sum_{b\in P_{\text{gt}}}\min_{a\in P_{\text{pre}}}\|a-b\|_{2}^{2}$ (1)
Similar to other self-supervised learning methods [27, 42], Point-MAE is typically trained on human-designed 3D shapes with semantic content (e.g., airplanes and chairs from ShapeNet [3]). In this work, we explore self-supervised learning with procedural 3D programs, referred to as Point-MAE-Zero. Here, “Zero” highlights that our 3D representations are learned entirely from procedurally generated 3D shapes, with zero human involvement in the generation process beyond the initial programming.
@@
We evaluate Point-MAE-Zero in various downsteam 3D understanding tasks, as shown in Fig. 1c. For classification tasks, we augment the pretrained transformer encoder with a three-layer MLP, which serves as the classification head.
For part segmentation tasks, We aggregate features from the 4th, 8th, and final layers of the Transformer encoder, and upsample the features to obtain the features for all 2048 input points. We include more implementation details in the supplementary material, and refer to Point-MAE [19] for more implementation details. For masked point cloud reconstruction, we use both pretrained transformer encoder and decoder, with any architecture modification.
Implementation details.
We closely follows Point-MAE [19] for the training of Point-MAE-Zero. Specifically, a input point cloud with $p=1024$ points is divided into $n=64$ patches, with each patch containing a constant number of points using $k=32$ in the KNN algorithm. The autoencoder comprises an encoder with 12 Transformer blocks and a decoder with 4 Transformer blocks. Each block includes 384 hidden dimensions, 6 attention heads. At the pretraining stage, we randomly sample 1024 points and apply standard random scaling and random translation for data augmentation. We pre-train our model for 300 epochs, using the AdamW optimizer [17] and cosine learning rate decay [16], with an initial learning rate of 0.001, a weight decay of 0.05, and a batch size of 128.
@@ 4 Experiments
In this section, we present a comprehensive evaluation of 3D shape representations pretrained with procedural 3D programs across various downstream tasks, including object classification, part segmentation, and masked point cloud completion (Sec. 4.1 – 4.3). We further provide an in-depth analysis of model behavior and ablation studies (Sec. 4.4). For each downstream task, we report the performance of relevant existing methods as a reference (primarily from Point-MAE [19]) and focus on comparisons with Point-MAE pretrained on ShapeNet, as well as models trained from scratch. Specifically, we evaluate the following three pretraining strategies:
  • Scratch: The network parameters are randomly initialized, with no pretraining.
  • Point-MAE-SN: The network is pretrained with Point-MAE on the ShapeNet (SN) [3] training split, containing 41,952 clean 3D models. We utilize the officially released pretrained Point-MAE model for this setting.
  • Point-MAE-Zero: The network is pretrained with Point-MAE on procedurally generated 3D models, using 150K synthetic models by default.

### v2

@@ Abstract
Remarkably, despite lacking semantic content, the 3D representations learned from the procedurally generated 3D shapes perform on par with state-of-the-art representations learned from semantically recognizable 3D models (e.g., airplanes) across various downstream 3D tasks, including shape classification, part segmentation, and masked point cloud completion. We provide a detailed analysis on factors that make a good 3D procedural programs. Extensive experiments further suggest that current self-supervised learning methods on point clouds do not rely on semantics of 3D shapes, shedding light on the nature of 3D representations learned.
@@ 1 Introduction
Self-supervised learning (SSL) aims at learning representations from unlabeled data that can transfer effectively to various downstream tasks. Inspired by the success of SSL in language [10] and image representation learning  [14, 15], SSL for 3D point cloud understanding has gained considerable interest [37, 55, 26]. Recently, Point-MAE [26] and its follow-ups [59, 61, 41] exploit the masked autoencoding scheme for 3D point cloud representation learning, showing substantial improvements in various 3D shape understanding tasks (e.g., shape classification, part segmentation, and scene instance segmentation).
However, unlike language and image data, which are abundantly available, 3D assets are less accessible due to the expertise required for creating 3D shapes using specialized software (e.g., Blender) or professional 3D scanners (e.g. LiDAR sensors). This scarcity of 3D shapes, known as data desert [11], limits the scalability of existing representation learning methods. Recent efforts have sought to expand 3D point cloud datasets at both the object level [8, 9] and scene level [3, 12, 53]. However, challenges unique to 3D data collection — such as copyright issues, format diversity, and scalability — remain unresolved.
Figure 1: Self-supervised learning from (a) procedurally generated 3D shapes [45, 49, 50, 17] performs comparably to learning from (b) ShapeNet models that are semantically meaningful [4] across various downstream 3D understanding tasks. Both outperforms training from scratch significantly. In (c), the x-axis represents various tasks and benchmarks: ModelNet40 [22] and three variants of ScanObjectNN [34] for shape classification, and ShapeNetPart [54] for part segmentation.
To address these challenges, we explore learning point cloud representations from solely synthetic data generated via procedural 3D programs [45, 49, 50, 17] — examples are shown in Fig. 1a. Our data generation pipeline begins with sampling shapes from a set of simple 3D primitives (e.g., cubes, cylinders, and spheres). These primitives undergo affine transformations (e.g., scaling, translation, and rotation) and are combined to create diverse geometries. We then augment the composed shapes using predefined operations (e.g., Boolean operations) to further enhance topological diversity, and uniformly sample 3D surface points from these shapes for point cloud representation learning. We generate 150K synthetic 3D point clouds in  600 CPU hours, with a scalable pipeline capable of producing unlimited 3D shapes free of copyright concerns. Additionally, we generate 2K synthetic scenes, including room layouts and objects following the described steps, requiring  500 CPU hours. This leads us to two key questions:
@@
We investigate learning 3D representations exclusively from synthetic shapes generated via procedural programs, without relying on human-crafted data. To validate this approach, we conduct extensive benchmarks across representative self-supervised learning methods, including Point-MAE [26], Point-M2AE [59], PCP-MAE [61], and Masked Scene Contrast (MSC) [41]. Specifically, we evaluate 3D object SSL methods on tasks such as shape classification, part segmentation, and masked point cloud reconstruction, while 3D scene SSL methods are evaluated on semantic and instance segmentation tasks. Our main findings are:
  • Despite lacking semantics, SSLs from solely synthetic data achieve performance comparable to their ShapeNet-pretrained counterparts. We also validate the effectiveness of procedurally generated data in scene-level 3D understanding tasks (Fig. 1c, Tab. 1 - Tab. 5).
  • We provide detailed insights into the factors that influence the quality of procedurally generated 3D datasets. We observe that learning performance improves notably with greater geometric diversity and increased dataset size (Tab. 4 and Fig. 5).
  • Our in-depth analysis reveals structural similarities between 3D shape representations from models pretrained on synthetic 3D shapes and semantically meaningful 3D shapes, providing insights into the nature of the learned representations (t-SNE visualization in Fig. 7).
To our best knowledge, this is the first systematic large-scale study on SSLs from procedural 3D programs. Our work is inspired by and builds upon a recent study that successfully trained large 3D reconstruction models exclusively on procedurally generated shapes [45, 17]. Our exploration is also closely related to prior efforts that learn image representations from procedural programs [2, 1]. Concurrent to our work, Yu et al. [56] demonstrated that procedurally generated videos can perform as effectively as natural videos for self-supervised video representation learning. Furthermore, our study is orthogonal yet complementary to recent efforts in scaling up 3D shape datasets [8, 9, 53].
@@ 2 Related Work
Figure 2: Learning from procedural 3D programs. (a) Synthetic 3D point clouds are generated by sampling, compositing, and augmenting simple primitives using procedural 3D programs [45]. (b) We experiment with multiple state-of-the-art self-supervised learning frameworks for learning 3D representations from synthetic data. Here, we illustrate the pretraining pipeline using Point-MAE [55], naming this variant Point-MAE-Zero, where “Zero” emphasizes the absence of any human-made 3D shapes. (c) We evaluate the pretrained models across various 3D shape understanding tasks.
@@
3D Object Datasets. Large-scale datasets are essential in advancing 3D deep learning. Unlike 2D images or videos, building and annotating 3D models requires expertise with professional 3D software or scanning equipment, making the process more costly and time-intensive. Despite these challenges, significant efforts have been made to curate extensive 3D shape datasets [4, 43, 30, 32, 6, 12, 25, 39, 8]. For example, ShapeNet provides 3 million CAD models, with 51K of them being clean, high-quality models, which serves as the standard benchmark for training and evaluating models in the 3D object recognition community. A more significant limitation of ShapeNet is its strong bias towards rigid, man-made artifacts, reflecting the inherent bias of its source 3D model repositories. More recently, Objaverse-XL [9] expanded the 3D dataset to 10.2 million models, though scaling up has introduced challenges, including increased noise, format diversity, and unresolved copyright and legal issues. In contrast, we explore procedural 3D programs that generate 3D shapes from simple primitives. This synthetic approach can theoretically produce an unlimited number of 3D shapes without licensing concerns. Moreover, procedural 3D programs provide a principled approach to mitigating biases inherently present in manually curated datasets [1].
@@
Supervised Learning for 3D Point Clouds. Unlike 2D images, 3D point clouds are inherently unordered and have an irregular structure, leading to extensive research on specialized neural network architectures [27, 28, 38, 62, 40, 42, 48, 19]. For instance, PointNet [27] introduces permutation-invariant operators and pooling layers feature aggregation across 3D points. PointNet++ [28] builds on this by incorporating a hierarchical spatial structure to capture more localized geometric information. DGCNN [38] adopts a graph-based approach, constructing a graph from input point clouds and applying a graph CNN to aggregate features from unordered 3D points. More recent models, such as the Point Transformer [62, 40, 42], utilize a modified transformer architecture tailored specifically for 3D point cloud processing. In our work, we employ a standard transformer architecture, aligning with recent trends in self-supervised learning [15, 26, 55].
@@
Self-supervised Learning for Point Clouds. Self-supervised learning (SSL) for 3D representations aims to learn features that transfer well across diverse point cloud tasks, such as shape classification, object detection, and part segmentation. Recent SSL approaches for point clouds generally fall into two categories: contrastive learning [14, 46, 5] and masked autoencoding [37, 55, 26, 11, 29, 13, 60, 57, 58]. PointContrast [46] and DepthContrast [5] use an instance discrimination task [14] to learn 3D representations. OcCo [37] learns point cloud representations by reconstructing the original point cloud from occluded views, while IAE [51] trains an autoencoder to recover implicit features from point cloud inputs. Inspired by the success of masked autoencoding in vision [15] and language representation learning [10], Point-BERT [55] and Point-MAE [26] learn representations by predicting masked regions, using transformers as the underlying architecture. More works have been built upon Point-MAE: Point-M2AE [59] introduces a pyramid architecture for both the encoder and decoder to incorporate multi-scale and hierarchical information in 3D shapes. More recently, PCP-MAE [61] highlights the issue of centroid leakage, which trivializes the pretext task in the PointMAE framework, and proposes predicting centroids as an additional objective to strengthen masked point reconstruction. 3D scene SSLs directly takes in scene-centric point clouds is an exciting research direction [41, 16, 47, 36, 21]. In this work, we adopt PointMAE, Point-M2AE, PCP-MAE and MSC, as our primary self-supervised learning framework due to its state-of-the-art performance across various 3D understanding tasks. For ablation studies, we use Point-MAE as the baseline, as it serves as the foundational approach in this research direction and provides the most representative evaluation.
@@
Learning from Synthetic Data Synthetic data has become popular in computer vision, especially in scenarios where ground-truth annotations are difficult to obtain or where privacy and copyright issues arise. State-of-the-art performance in mid-level or 3D vision tasks is often achieved through training on synthetic data, including tasks like optical flow [33], depth estimation [52], dense tracking [18], relighting [49], novel view synthesis [50], and material estimation [20]. Procedurally generated synthetic data has also been explored for self-supervised representation learning in images [2, 1] and videos [56], and more recently for multi-view feed-forward 3D reconstruction [45]. Concurrent with our study, Yu et al. [56] investigates self-supervised video representation learning from procedurally generated images and videos. In this work, we explore self-supervised representation learning for point clouds, using synthetic 3D shapes generated by procedural 3D programs.
@@ 3.1 Procedural 3D programs
There is a line of work synthesizing procedural 3D shapes for vision tasks such as novel view synthesis [50], relighting [49], and material estimation [20]. Following recent methods [45] that use procedural 3D shapes for sparse-view reconstruction, we address self-supervised 3D representation learning from purely synthetic datasets. Fig. 2a illustrates our data pipeline:
  • (1) Randomly sample $\mathbf{K}$ primitive shapes (cubes, spheres, cylinders, cones, tori) and apply affine transformations to combine them;
  • (2) Apply geometric augmentations (e.g., boolean differences, wireframe conversions) to enrich shape diversity (see [45] for details);
  • (3) Uniformly sample $\mathbf{N}$ surface points per synthesized shape as inputs for representation learning (Fig. 2b).
We experiment with various shape-generation configurations, such as changing the number of sampled primitives and applying augmentations. By default, each dataset consists of 150K shapes with $\mathbf{N}=8192$ points each. Sec. 4 further analyzes the effects of dataset size and shape complexity on learned representations.
In order to generate procedural 3D scenes, we follow MegaSynth [17] to procedurally generate 2K synthetic 3D scenes. Specifically, we first generate a floor plan and generate procedural 3D shapes with the above pipeline and place procedural 3D shapes in the scene based on the generated floor plan. We include details on how we generate 2K procedural 3D scenes in the supplementary material.
3.2 Procedural Pretraining
@@
As depicted in Fig. 2b, we adopt Point-MAE [15], Point-M2AE [59], PCP-MAE [61] to train on procedurally generated 3D shapes. These methods rely on a masked autoencoding scheme [10, 15, 55], where the input point cloud is split into irregular patches and a large portion of them (60% by default) is randomly masked. A Transformer-based encoder–decoder network then attempts to reconstruct these masked patches, thereby learning 3D representations. Concretely, the reconstruction loss is computed as the $L_{2}$ Chamfer Distance between the predicted point patches $P_{\text{pre}}$ and the ground-truth patches $P_{\text{gt}}$:
$L=\frac{1}{|P_{\text{pre}}|}\sum_{a\in P_{\text{pre}}}\min_{b\in P_{\text{gt}}}\|a-b\|_{2}^{2}+\frac{1}{|P_{\text{gt}}|}\sum_{b\in P_{\text{gt}}}\min_{a\in P_{\text{pre}}}\|a-b\|_{2}^{2}.$ (1)
For scene-level SSLs, we adopt MSC [41], which combines masked auto-encoding and contrasive learning, to train on procedurally generated 3D scenes. While earlier self-supervised approaches [37, 55] often train on human-crafted 3D models (e.g., ShapeNet [4]), here we focus on purely synthetic data. We thus call add suffix -Zero to emphasize that it is trained exclusively on procedural 3D shapes with zero human-designed content. For clarity, we use the suffix -SN to indicate baselines trained on ShapeNet.
@@
We evaluate baselines on several 3D tasks, as summarized in Fig. 2c. For shape classification, we augment the pretrained Transformer encoder with a three-layer MLP classification head. For part segmentation, we aggregate features from the 4th, 8th, and final layers of the encoder, upsample them to all 2048 input points, and employ a segmentation head. For masked point cloud reconstruction, we use both the pretrained encoder and decoder with no architectural modifications. For scene-level methods, we use both instance segmentation and semantic segmentation finetuned from the pretrained SSLs with a linear prediction head. Detailed implementation settings are in the supplementary material.
@@ 4 Experiments
We present a comprehensive evaluation of 3D shape representations pretrained with procedural 3D programs across various downstream object-level and scene-level tasks, including object classification, part segmentation, and 3D scene understanding (Sec. 4.1 – 4.4). We further provide an in-depth analysis of model behavior and ablation studies (Sec. 4.5). For each downstream task, we report the performance of relevant existing methods as a reference and focus on comparisons with SSLs pretrained on manually-curated 3D datasets, as well as models trained from scratch. Specifically, for object-level 3D understanding tasks, we evaluate the following three pretraining strategies: (1) Scratch: All network parameters are randomly initialized, with no pretraining. (2) ShapeNet Pretrained (SN): Pretrained on 41,952 models in the ShapeNet [4] training split, relying on the officially released weights. (3) Procedural 3D Programs Pretrained (Zero): Pretrained on 150K procedurally generated 3D models, using no human-crafted shapes. For scene-level tasks, we compare SSL models pretrained on ScanNet [7] and procedurally generated 3D scenes.